# Reduced-Order Modeling of Unsteady Fluid Flow Using Neural Network Ensembles

### v1

@@ Abstract
Keywords model reduction  $\cdot$ deep learning  $\cdot$ non-intrusive  $\cdot$ machine learning  $\cdot$ ensemble learning  $\cdot$ computational fluid dynamics
@@ 1 Introduction
Employing reduced-order models (ROMs) is a prevalent strategy to mitigate this high computational cost. By utilizing a limited set of training data from computed high-fidelity simulations, ROMs construct a low-dimensional surrogate model that can be evaluated in real-time and provide accurate full-order model approximations at unseen designs within the parameter space covered by the training data. ROMs significantly reduce the degrees of freedom of the full-order model through the use of a low-dimensional embedding that can be effectively mapped back to the full-order state. ROMs consist of two stages: an offline stage, which is computationally intensive, involving the computation of high-fidelity solutions by solving the full-order model (FOM) to generate data snapshots and train the low-dimensional surrogate model, and an online stage, where the model is used to approximate solutions at desired points.
@@
Most ROMs utilize the proper orthogonal decomposition [1], a linear method that uses the singular value decomposition (SVD) to obtain a low-rank subspace consisting of a number of linearly independent basis vectors. A linear combination of these basis vectors is computed using a set of expansion coefficients to approximate full-order states within the solution space. Although POD is widely used, it encounters difficulties when applied to highly nonlinear problems, often requiring a large number of basis vectors to provide reasonable accuracy [2].
@@
More recently, machine learning and artificial intelligence (AI) methods have been used to provide nonlinear mappings between the low-dimensional embedding and high-fidelity solution space. In particular, deep learning [3] approaches have been used to develop ROMs that provide efficient nonlinear manifolds of physical systems. Convolutional autoencoders (CAEs), a type of artificial neural network, have been used to build ROMs and shown to outperform POD-based methods  [4, 5]. Artificial neural networks that use convolutional layers efficiently learn structures and patterns present in spatially distributed data, including the solutions to PDEs. Autoencoders consist of two individual neural networks: an encoder, which maps high-dimensional inputs to a low-dimensional latent space, and a decoder, which maps the low-dimensional latent space to a reconstruction of the high-dimensional input.
@@
ROMs are categorized as either non-intrusive or projection-based; non-intrusive ROMs use a fully data-driven approach, while projection-based ROMs incorporate the governing equations to solve a low-dimensional version of the full-order model. While projection-based ROMs can provide better accuracy and more physically consistent results, they incur a larger computational cost and often exhibit stability issues [6] that inhibit convergence. Additionally, they are generally not portable between different solvers.
@@
ROMs are developed for both steady and unsteady physics problems. The first approach involves computing a single point estimate of the low-dimensional embedding, while the latter involves making predictions over a prescribed time horizon. Unsteady non-intrusive ROMs combine either POD or CAE for spatial reconstruction of full-order states with a model used to make time-series predictions of the low-dimensional embeddings. Deep learning methods are popular for this as well, including long short-term memory (LSTM) neural networks [7] and transformer neural networks [8]. Both LSTMs and transformers are powerful models for handling data that are sequential in nature such as time-series data. Transformers utilize an architecture that is much more complex than the one LSTMs use, which leads to higher computational costs for both training and inference.
@@ 2.1 Proper Orthogonal Decomposition
A subspace $\bm{\mathcal{V}}$ associated with $\bm{X}$ exists such that $\bm{\mathcal{V}}$ = span$(\bm{X})$. It it assumed that $\bm{\mathcal{V}}$ provides a good approximation of the solutions for $\bm{\mu}$ within the parameter space if there is a sufficient variety of training data in $\bm{X}$. A rank $k$ set of orthonormal basis vectors $[{\bm{\psi}}^{1},{\bm{\psi}}^{2},\cdots,{\bm{\psi}}^{k}]\in\mathbb{R}^{N}$, where $k\ll N$, is associated with $\bm{\mathcal{V}}$ such that each solution $\bm{x}^{i}$ in $\bm{X}$ can be reconstructed as a linear combination of the basis vectors
@@
where $\bm{a}^{i}$ are the low-dimensional expansion coefficients. The truncated singular value decomposition (SVD) of $\bm{X}$ decomposes the matrix into two unitary matrices $\bm{U}\in\mathbb{R}^{N\times n}$ and $\bm{V}\in\mathbb{R}^{n\times n}$, and a diagonal matrix $\bm{\Sigma}\in\mathbb{R}^{n\times n}$ such that
@@ 2.3 Long Short-Term memory (LSTM) Neural Networks
A unique set of weights and biases belongs to each gate or cell state. When using LSTMs for time-series prediction, an autoregressive prediction is used, often referred to as the sliding window approach. For a given input sequence containing multivariate data for $w$ timesteps, the next step in the sequence is predicted. When making the next prediction, the current prediction is incorporated into the input sequence by removing the first element and shifting the rest to the prior position. Eventually, the model inputs will consist of only previously computed predictions, which makes model performance sensitive to error propagation. The input sequence length, or window size $w$, is an important hyperparameter to consider when training LSTMs. Although LSTMs are designed to effectively learn long-range dependencies, using an input sequence length that is too long can lead to the inclusion of outdated and irrelevant information. On the other hand, using an input sequence length that is too small can ignore important long-range information. A diagram of a single-layer LSTM making a prediction one timestep ahead is shown in Figure 2.
@@ 2.4 Ensemble Learning
Bagging reduces the issue of error propagation by averaging the errors of multiple weak learners, which will become increasingly small as the number of learners grows, leading to much lower variance. Bagging also offers a considerable gain in robustness and generalization by leveraging the diversity of the weak learners to increase the capability of handling varying patterns outside of the training data.
@@ 2.5 CAE-eLSTM ROM
This section describes the non-intrusive ROM framework combining convolutional autoencoders for spatial reconstruction and LSTM ensembles utilizing bagging for temporal prediction, which we will refer to as a CAE-eLSTM ROM. Solutions to the full-order model for designs $\bm{\mu}$ in the training set of design parameters $\bm{\mathcal{U}}_{\text{train}}\in\mathbb{R}^{n\times p}$ are obtained during the offline stage and are assembled into a matrix $\bm{X}$. A convolutional autoencoder with latent dimension $k$ is trained on $\bm{X}$ for a sufficient number of epochs such that inputs are accurately reconstructed. Next, the expansion coefficients for training data $\bm{A}_{\text{train}}$ are computed using the encoder $g_{\text{enc}}$. Sequences of length $w$ are then generated from $\bm{A}_{\text{train}}$ and $m$ bagged LSTMs are trained on individual data subsets chosen randomly through sampling with replacement.
The online stage involves executing the ROM to compute approximate solutions at a point $\bm{\mu}^{*}$, where the full-order model is initially run for $T_{i}$ timesteps. An initial sequence of latent variables of length $w$ is required as an input to the ROM. This means that the FOM must be run for at least $w$ timesteps, and potentially longer depending on the how useful initial simulation data is for the ROM. The latent variables for the rest of the simulation are calculated autoregressively using the bagged LSTMs by taking an average of the individual predictions. The online and offline stages are outlined in Algorithm 1. When using a ROM with a single LSTM as in previous works [12, 13], the LSTM is trained on the entire time-series dataset and is used alone for time-series prediction.
@@
2:     Run FOM for $T_{i}$ timesteps at $\bm{\mu}^{*}$ and compute the expansion coefficients $\bm{A}^{*}_{T_{i}}=g_{\text{enc}}\left(\bm{X}_{T_{i}}\right)$.
@@
7:     Predict full-order snapshots $\bm{\tilde{X}}^{*}=g_{\text{dec}}\left(\bm{{A}}^{*}\right)$
@@ 3 Results
The two test cases used to demonstrate the performance of the CAE-eLSTM ROM are a lid-driven cavity and the flow over a cylinder. Both cases use two-dimensional computational domains. The lid-driven cavity case consists of three design parameters controlling the geometry of the domain, while the cylinder case consists of two design parameters controlling geometric and physical properties of the simulation. Latin hypercube sampling (LHS) [25], a popular statistical method for generating samples from a multi-dimensional parameter space is used to generate sets of design parameters that are split into training and test sets. LHS aims to maximize the distance and minimize the correlation amongst samples produced by using uniform sampling. The ROM is used to predict the vertical and horizontal components of the velocity. The training data for each velocity component is scaled to a range [0,1] using min-max scaling before input into the CAE. Similarly, the CAE latent variables are also scaled using min-max scaling before being used for the LSTM. Data normalization improves performance and allows the optimizer to learn optimal network parameters at a much faster rate [26]. Both the CAE and LSTM are trained using the mean squared error loss function. An NVIDIA DGX system, consisting of eight NVIDIA A100 GPUs, is used to train the CAE and LSTMs, run the cylinder simulations, and for ROM inference. The lid-driven cavity simulations are run on a local workstation using a single CPU core.
The performance of the ensemble ROM is tested against a CAE-LSTM ROM that uses a single LSTM model. Using five different random initializations of weights and biases prescribed by different seeds, the time-series of latent variables are visually compared in addition to performance metrics averaged over the seeds. When training multiple LSTMs on bootstrapped data sets for a single seed, the same initial weights and biases are used. $N_{s}=5$ seeds are used to initialize the LSTM model, with $s=$[1, 2, 3, 4, 5]. For a single test point $\bm{\mu}^{*}$ and seed $s$, an error metric of interest is the relative $L^{2}$ error in a field component averaged over the prediction time horizon,
@@
A lower standard deviation in the error term indicates that the model performance does not vary much with the initial weights and biases.
@@ 3.1 Lid-Driven Cavity
100 sets of design parameters are generated and randomly split into 90 training samples and 10 test points, which are given in Table 1. The CAE-eLSTM ROM uses $m$ = 64 bagged LSTMs and a window size $w=20$. These parameters were chosen through a trial-and-error process with the goal of maximizing predictive accuracy while trying to keep the computational cost of training the LSTM ensemble relatively low. The CAE latent dimension is set to $k=4$; below this value, the CAE reconstruction errors were found to be worse, and increasing $k$ offered no improvement. The LSTM architecture consists of two hidden layers consisting of 50 neurons each and a dropout [28] rate of 0.1. The output layer contains a sigmoid activation function so the outputs are constrained to a range of $[0,1]$. The CAE architecture is given in Appendix B. The Adam optimizer [29] is used to train both the CAE and LSTM, with an initial learning rate of $\eta=5\times 10^{-4}$ and weight decay of $\lambda=1\times 10^{-6}$. Dropout and weight decay are used to prevent overfitting. The CAE is trained for a total of 200 epochs, while an individual LSTM is trained for 250 epochs. At the test points, the FOM is simulated for $T_{i}=0.75$ seconds (15 snapshots), or 15% of the total time horizon.
@@
Table 1: Test case design parameters for the lid-driven cavity problem.
Figure 7 shows the trajectories of the latent variables computed using ensemble and single LSTMs at the test point $\bm{\mu^{*}}=[1.299,1.689,-0.0367]$ using five different seeds which control the initial weights and biases of the LSTM. It is shown that for all of the latent variables, using LSTM ensembles leads to higher prediction accuracy and significantly lower variance. When using a single LSTM, the different predictions quickly diverge from each other and the ground truth. In contrast, the ensemble model is much less sensitive to the initial seed, and the predicted trajectories do not differ much. The second latent variable exhibits diverging trajectories when using the ensemble model, but the effect is much less pronounced than when using a single LSTM. Figure 6 shows contours of $u$ and $v$ at the test point as well as absolute ROM errors averaged over the last 10% of the simulation ($t\in$ [4.5, 5] seconds) for a single seed. The errors are considerably larger when using a single LSTM throughout the computational domain for both $u$ and $v$. Table 2 lists the seed-averaged relative errors $\bar{\epsilon}$ in $u$ and $v$ for the test points. The ensemble method offers better performance in predicting both fields at all of the test points, usually by wide margins. Table 3 lists the standard deviation of these errors; at all of the test points, the standard deviation is also smaller, showing that using the ensemble method leads to much greater stability and reliability in predictions. Table 4 lists the computational costs associated with CFD simulation and ROM inference. The given CFD wall time is the portion of the simulation over the prediction time horizon; using the ROM for inference offers a speed-up of approximately 9.4x over CFD.
@@
Table 2: Average relative errors in $u$ and $v$ for the lid-driven cavity case
@@
Table 3: Standard deviation in relative errors of $u$ and $v$ for the lid-driven cavity case.
@@
Table 4: Computational costs associated with the lid-driven cavity problem.
@@
The next test case involves two-dimensional incompressible, unsteady, laminar flow over a cylinder between two solid walls. Eventually, the lid-driven cavity flow from the previous test case reaches steady-state and does not exhibit long-term transient behavior that is commonly found in fluid dynamics problems. Laminar flow over a cylinder is a well-studied problem in fluid dynamics, with both experimental and computational results present in the literature [30, 31]. Unsteady cylinder flow is characterized by the presence of vortices that separate from the surface and form in the wake. This distinctive pattern is known as the Von Kármán vortex street, where alternating vortices of a regular pattern are shed downstream of the cylinder. The unsteady Navier-Stokes equations found in equations 14 and 15 are solved using XLB [32], a Lattice Boltzmann method [33] library utilizing the JAX framework [34] available for Python, which allows for effective scaling across multiple CPUs, GPUs, and distributed multi-GPU systems. The cited works can be referred to for an overview of the Lattice Boltzmann method and its implementation in XLB.
@@
The diameter $d$ is set to an integer quantity by rounding to the nearest whole number. The wake structure behind the cylinder undergoes instabilities [35, 36] at a critical reynolds number of approximately $Re_{c}=180$, where the vortices transition to becoming turbulent. As a result, the prescribed range given for $Re$ in the parameter space includes a variety of physical regimes. Additionally, both $Re$ and $d$ control the size and periodicity of the shed vortices, making this a difficult prediction problem for ROMs. The freestream velocity in the $x$-direction is set to $u_{\infty}=0.001$ in non-dimensional units. The simulation is run for $T=750000$ timesteps with data being saved every 1000 steps, leading to 750 snapshots for a single simulation. The initial condition for a simulation of a given diameter $d$ is the solution to steady flow at $Re=20$.
50 sets of design parameters are generated and split into 45 training samples and 5 test points, shown in Figure 8. Table 5 also lists the test point design parameters. The ROM uses $m=96$ bagged LSTMs and a window size of $w$ = 30, which were again chosen through a trial-and-error process to maximize accuracy while keeping the number of weak learners to a minimum. The CAE latent dimension is set to $k=10$. Similar to the lid-driven cavity case, values below $k=10$ resulted in higher reconstruction errors and increasing the latent dimension further offers no improvement. A bidirectional LSTM architecture [37] is used. Bidirectional recurrent neural networks process sequential data in both forward and backward directions, allowing the model to learn both past and future context. The network consists of three hidden layers with 36 neurons each and a dropout rate of 0.15. The output layer again contains a sigmoid activation function, and the Adam optimizer with the same initial learning rate $\eta=5\times 10^{-4}$ and weight decay of $\lambda=1\times 10^{-6}$ is used for both the LSTM and CAE. Again, the CAE is trained for 200 epochs while an individual LSTM is trained for 250 epochs. The CAE architecture is given in Appendix B. At the test points, the full-order model is run for $T_{i}=300000$ seconds (300 snapshots), or 40% of the total simulation time. This value is required to be high as the flow exhibits highly oscillatory behavior initially, leading to very noisy latent variables that cannot be used for model training. As a result, latent variable sequences are generated starting at the 200th snapshot, and $\bm{A}_{\text{train}}$ does not contain time-series data of latent variables before this point.
@@
Table 5: Test case design parameters for the 2D cylinder case.

### v2

@@ 1 Introduction
Employing reduced-order models (ROMs) is a prevalent strategy to mitigate this high computational cost [1]. By utilizing a limited set of training data from computed high-fidelity simulations, ROMs construct a low-dimensional surrogate model that can be evaluated in real-time and provide accurate full-order model (FOM) approximations at unseen designs within the parameter space covered by the training data. ROMs significantly reduce the degrees of freedom of the FOM through the use of a low-dimensional embedding that can be effectively mapped back to the full-order state, leading to a substantial decrease in computational cost. ROMs consist of two stages: an offline stage, which is computationally intensive, involving the computation of high-fidelity solutions by solving the FOM to generate data snapshots and train the low-dimensional surrogate model, and an online stage, where the model is used to approximate solutions at desired points. Running the FOM requires numerically solving the governing equations over the computational domain, which comes at a much larger computational cost than the ROM, which represents the FOM using a small number of variables.
@@
Most ROMs utilize the proper orthogonal decomposition [2], a linear method that uses the singular value decomposition (SVD) to obtain a low-rank subspace consisting of a number of linearly independent basis vectors. A linear combination of these basis vectors is computed using a set of expansion coefficients to approximate full-order states within the solution space. The POD basis vectors are interpretable, allowing for visualization of dominant physical features. Although POD is widely used, it encounters difficulties when applied to highly nonlinear problems, often requiring a large number of basis vectors to provide reasonable accuracy [3].
@@
More recently, machine learning and artificial intelligence (AI) methods have been used to provide nonlinear mappings between the low-dimensional embedding and high-fidelity solution space. In particular, deep learning [4] approaches have been used to develop ROMs that provide efficient nonlinear manifolds of physical systems. Convolutional autoencoders (CAEs), a type of artificial neural network, have been used to build ROMs and shown to outperform POD-based methods  [5, 6]. Artificial neural networks that use convolutional layers efficiently learn structures and patterns present in spatially distributed data, including the solutions to PDEs. Autoencoders consist of two individual neural networks: an encoder, which maps high-dimensional inputs to a low-dimensional latent space, and a decoder, which maps the low-dimensional latent space to a reconstruction of the high-dimensional input. Although autoencoders can effectively learn nonlinear relationships, they lack interpretability. Recent works have used variational autoencoders [7, 8] to incorporate interpretable nonlinear modes into ROMs, although this results in lower reconstruction accuracy when compared to vanilla autoencoders. Another approach for simulating systems governed by PDEs at a reduced cost is the use of neural operators [9, 10], which implicitly learn the governing equations using deep neural networks. However, this approach typically requires very large amounts of training data to cover parameter spaces effectively, limiting their use in design optimization.
@@
ROMs are categorized as either non-intrusive or projection-based; non-intrusive ROMs use a fully data-driven approach, while projection-based ROMs incorporate the governing equations to solve a low-dimensional version of the FOM. While projection-based ROMs can provide better accuracy and more physically consistent results, they incur a larger computational cost and often exhibit stability issues [11] that inhibit convergence. Additionally, they are generally not portable between different solvers.
@@
ROMs are developed for both steady and unsteady physics problems. The first approach involves computing a single point estimate of the low-dimensional embedding, while the latter involves making predictions over a prescribed time horizon. Unsteady non-intrusive ROMs combine either POD or CAE for spatial reconstruction of full-order states with a model used to make time-series predictions of the low-dimensional embeddings. Deep learning methods are popular for this as well, including long short-term memory (LSTM) neural networks [12] and transformer neural networks [8]. Both LSTMs and transformers are powerful models for handling data that are sequential in nature such as time-series data. Transformers utilize an architecture that is much more complex than the one LSTMs use, which can lead to better performance for complex time-series problems and other tasks such as developing large language models [13]. However, this comes at an increased cost for both training and inference when compared to LSTMs, which use a significantly lower number of parameters. The use of LSTMs for time-series forecasting is well-established [14, 15], making them a popular choice.
@@ 2.1 Proper Orthogonal Decomposition
A subspace $\bm{\mathcal{V}}$ associated with $\bm{X}$ exists such that $\bm{\mathcal{V}}$ = span$(\bm{X})$. It assumed that $\bm{\mathcal{V}}$ provides a good approximation of the solutions for $\bm{\mu}$ within the parameter space if there is a sufficient variety of training data in $\bm{X}$. A rank $k$ set of orthonormal basis vectors $[{\bm{\psi}}^{1},{\bm{\psi}}^{2},\cdots,{\bm{\psi}}^{k}]\in\mathbb{R}^{N}$, where $k\ll N$, is associated with $\bm{\mathcal{V}}$ such that each solution $\bm{x}^{i}$ in $\bm{X}$ can be reconstructed as a linear combination of the basis vectors
@@
where $\bm{a}^{i}$ are the low-dimensional expansion coefficients. The truncated singular value decomposition (SVD) of $\bm{X}$ decomposes the matrix into two orthonormal matrices $\bm{U}\in\mathbb{R}^{N\times n}$ and $\bm{V}\in\mathbb{R}^{n\times n}$, and a diagonal matrix $\bm{\Sigma}\in\mathbb{R}^{n\times n}$ such that
@@ 2.3 Long Short-Term memory (LSTM) Neural Networks
A unique set of weights and biases belongs to each gate or cell state. When using LSTMs for time-series prediction, an autoregressive prediction is used, often referred to as the sliding window approach. For a given input sequence containing multivariate data for $w$ timesteps, the next step in the sequence is predicted. When making the next prediction, the current prediction is incorporated into the input sequence by removing the first element and shifting the rest to the prior position. Eventually, the model inputs will consist of only previously computed predictions, which makes model performance sensitive to error propagation. The input sequence length, or window size $w$, is an important hyperparameter to consider when training LSTMs. Although LSTMs are designed to effectively learn long-range dependencies, using an input sequence length that is too long can lead to the inclusion of outdated and irrelevant information. On the other hand, using an input sequence length that is too small can ignore important long-range information. The optimal choice is highly problem-dependent; similar problems found in the literature can be consulted, or methods such as cross-validation can be used to find an optimal value. A diagram of a single-layer LSTM making a prediction one timestep ahead is shown in Figure 2.
@@ 2.4 Ensemble Learning
Bagging reduces the issue of error propagation by averaging the errors of multiple weak learners, which will become increasingly small as the number of learners grows, leading to much lower variance. Bagging also offers a considerable gain in robustness and generalization by leveraging the diversity of the weak learners to increase the capability of handling varying patterns outside of the training data. As an increasing number of weak learners are added, the reduction in variance plateaus, leading to diminishing returns in performance gains. After a certain point, the computational cost of adding more weak learners outweighs the marginal improvement in predictive performance. As the complexity of the problem increases, the number of weak learners required to make accurate predictions also generally increases. While theoretical considerations can be made to analyze the trade-offs between bias, variance, and covariance to determine the optimal number of weak learners, this number is difficult to choose a priori. A reasonable choice is based on both the problem’s complexity and similar problems found in the literature, which are not yet present in our case and are established in this work.
@@ 2.5 CAE-eLSTM ROM
This section describes the non-intrusive ROM framework combining convolutional autoencoders for spatial reconstruction and LSTM ensembles utilizing bagging for temporal prediction, which we will refer to as a CAE-eLSTM ROM. The offline stage is computationally intensive, requiring multiple solves of the FOM in addition to training multiple neural networks. First, solutions to the FOM for designs $\bm{\mu}$ in the training set of design parameters $\bm{\mathcal{U}}_{\text{train}}\in\mathbb{R}^{n\times p}$ are obtained for all timesteps and are assembled into a matrix $\bm{X}$. Next, a convolutional autoencoder with latent dimension $k$ is trained on $\bm{X}$ for a sufficient number of epochs such that inputs are accurately reconstructed. Then, the expansion coefficients for the training data $\bm{A}_{\text{train}}$ are computed by feeding the full-order states from $\bm{X}$ through the encoder $g_{\text{enc}}$. Sequences of length $w$ are then generated from $\bm{A}_{\text{train}}$ and $m$ LSTMs using the same set of initial weights and biases are trained on individual data subsets chosen randomly through sampling with replacement. PyTorch [32], an open source deep learning framework available for Python, is used to implement the autoencoder and LSTM using default weight and bias initialization settings.
A schematic describing the online stage of the ROM is shown in Figure 4. The online stage involves executing the ROM to compute approximate solutions at a point $\bm{\mu}^{*}$. First, the FOM is run for $T_{i}$ timesteps, as an initial sequence of latent variables is required for the LSTM ensemble. The computed full-order states are fed through the encoder to obtain the initial sequence of latent variables $\bm{A}^{*}_{T_{i}}\in\mathbb{R}^{k\times w}$. The FOM must be run for at least $T_{i}\geq w$ timesteps, and potentially longer depending on the how useful simulation data from the initial timesteps are for the ROM. The latent variables for the rest of the prediction horizon are calculated autoregressively using the LSTM ensemble by taking an average of the predictions from the individual weak learners. After the predicted latent variables are found for each timestep $t$, they are fed through the decoder to obtain approximate full-order fields. The online and offline stages are outlined in Algorithm 1.
When using a ROM with a single LSTM as in previous works [19, 20], the LSTM is trained on the entire time-series dataset and is used alone for time-series prediction. When compared to these works, the main difference in our presented ROM framework is the use of an ensemble model for time-series prediction. Although this results in an increased offline training cost, using an ensemble allows for better accuracy and stability over long prediction horizons. While the use of an ensemble LSTM is simple, it can greatly improve ROM performance without requiring additional training data.
@@
2:     Run FOM for $T_{i}$ timesteps at $\bm{\mu}^{*}$ and compute the expansion coefficients $\bm{A}^{*}_{T_{i}}=g_{\text{enc}}\left([\text{X}^{*}_{T_{1}},\text{X}^{*}_{T_{2}},\cdots\text{X}^{*}_{T_{i}}]\right)$.
@@
7:     Predict full-order fields $\bm{\tilde{X}}^{*}=g_{\text{dec}}\left(\bm{{A}}^{*}\right)$
@@ 3 Results
The two test cases used to demonstrate the performance of the CAE-eLSTM ROM are a lid-driven cavity and the flow over a cylinder. Both cases use two-dimensional computational domains. The lid-driven cavity case consists of three design parameters controlling the geometry of the domain, while the cylinder case consists of two design parameters controlling geometric and physical properties of the simulation. Different CFD solvers are used for each case; this is done to illustrate the portability of the ROM and allow for the use of a structured grid for the cylinder case. Latin hypercube sampling (LHS) [33], a popular statistical method for generating samples from a multi-dimensional parameter space is used to generate sets of design parameters that are split into training and test sets. LHS aims to maximize the distance and minimize the correlation amongst samples produced by using uniform sampling. The ROM is used to predict the vertical and horizontal components of the velocity. The training data for each velocity component is scaled to a range [0,1] using min-max scaling before input into the CAE. Similarly, the CAE latent variables are also scaled using min-max scaling before being used for the LSTM. Data normalization improves performance and allows the optimizer to learn optimal network parameters at a much faster rate [34]. Both the CAE and LSTM are trained using the mean squared error loss function. An NVIDIA DGX system, consisting of eight NVIDIA A100 GPUs, is used to train the CAE and LSTMs, run the cylinder simulations, and for ROM inference. The lid-driven cavity simulations are run on a local workstation using a single CPU core.
The performance of the ensemble ROM is tested against a CAE-LSTM ROM that uses a single LSTM model. Using five different random initializations of the weights and biases prescribed by different seeds, the time-series of latent variables are visually compared in addition to performance metrics averaged over the seeds. When training multiple LSTMs on bootstrapped data sets for a single seed $s$, the same initial weights and biases are used for each individual weak learner. Evaluating the ROM’s performance using different seeds allows for testing its sensitivity to the initial weights and biases. $N_{s}=5$ different seeds are used to initialize the LSTM model, with $s=$[1, 2, 3, 4, 5]. For a single test point $\bm{\mu}^{*}$ and seed $s$, an error metric of interest is the relative $L^{2}$ error in a field component averaged over the prediction time horizon,
@@
A lower standard deviation in the error term indicates that the model performance over the prediction horizon does not vary much with the initial weights and biases.
@@ 3.1 Lid-Driven Cavity
100 sets of design parameters are generated and randomly split into 90 training samples and 10 test points, which are given in Table I. The CAE-eLSTM ROM uses $m$ = 64 bagged LSTMs and a window size $w=20$. These parameters were chosen through a trial-and-error process with the goal of maximizing predictive accuracy while trying to keep the computational cost of training the LSTM ensemble relatively low. This was done by directly evaluating the performance on the test set, while in practice methods like cross-validation would be used. Initial guesses for the window size were based on similar problems found in the literature. An initial number of approximately 50 weak learners was used, which was increased in increments of 8 until the performance gains plateaued. The CAE latent dimension is set to $k=4$; below this value, the CAE reconstruction errors were found to be worse, and increasing $k$ offered no improvement. The LSTM architecture consists of two hidden layers consisting of 50 neurons each and a dropout [36] rate of 0.1. The output layer contains a sigmoid activation function so the outputs are constrained to a range of $[0,1]$. The CAE architecture is given in Appendix B. The Adam optimizer [37] is used to train both the CAE and LSTM, with an initial learning rate of $\eta=5\times 10^{-4}$ and weight decay of $\lambda=1\times 10^{-6}$. Dropout and weight decay are used to prevent overfitting. The CAE is trained for a total of 200 epochs, while an individual LSTM is trained for 250 epochs. At the test points, the FOM is simulated for $T_{i}=0.75$ seconds (30 snapshots), or 15% of the total time horizon. The final 20 of these snapshots are used as the initial input sequence for the LSTM ensemble.
@@
Table I: Test case design parameters for the lid-driven cavity problem.
Figure 7 shows the trajectories of the latent variables computed using ensemble and single LSTMs at the test point $\bm{\mu^{*}}=[1.299,1.689,-0.0367]$ using five different seeds which control the initial weights and biases of the LSTM. It is shown that for all of the latent variables, using LSTM ensembles leads to higher prediction accuracy and significantly lower variance. When using a single LSTM, the different predictions quickly diverge from each other and the ground truth. Due to their high capacity to learn, neural networks often exhibit high variance. Since neural networks contain a large number of parameters, the effects of error propagation between different trained models are heightened due to relatively large differences in autoregressive predictions being fed back into the model. In contrast, the ensemble model is much less sensitive to the initial seed, and the predicted trajectories do not differ much. The second latent variable exhibits diverging trajectories when using the ensemble model, but the effect is much less pronounced than when using a single LSTM. Figure 8 shows contours of $u$ and $v$ at the test point as well as absolute ROM errors averaged over the last 10% of the simulation ($t\in$ [4.5, 5] seconds) for a single seed. The errors are considerably larger when using a single LSTM throughout the computational domain for both $u$ and $v$. Table II lists the seed-averaged relative errors $\bar{\epsilon}$ in $u$ and $v$ for the test points. The ensemble method offers better performance in predicting both fields at all of the test points, usually by wide margins. Table III lists the standard deviation of these errors; at all of the test points, the standard deviation is also smaller. As there is a lower spread amongst the different seeds in the time-averaged error metrics, this shows that using the ensemble method leads to much greater stability and reliability in predictions. Table IV lists the computational costs associated with neural network training, CFD simulation, and ROM inference. The given CFD wall time is the portion of the simulation over the prediction time horizon; using the ROM for inference over this period of time offers a speed-up of approximately 9.4x over CFD. The ROM inference costs include autoregressive prediction using the ensemble LSTM and obtaining full-order approximations using the decoder. The CAE is much more expensive to train than a single LSTM as it has a larger number of trainable parameters. While a multi-GPU architecture was used to train the LSTMs in parallel, the cost to train a single LSTM is given.
@@
Table II: Average relative errors in $u$ and $v$ for the lid-driven cavity case. Boldface denotes the lowest values.
@@
Table III: Standard deviation in relative errors of $u$ and $v$ for the lid-driven cavity case. Boldface denotes the lowest values.
@@
Table IV: Computational costs associated with the lid-driven cavity problem.
@@
The next test case involves two-dimensional incompressible, unsteady, laminar flow over a cylinder between two solid walls. Eventually, the lid-driven cavity flow from the previous test case reaches steady-state and does not exhibit long-term transient behavior that is commonly found in fluid dynamics problems. Laminar flow over a cylinder is a well-studied problem in fluid dynamics, with both experimental and computational results present in the literature [38, 39]. Unsteady cylinder flow is characterized by the presence of vortices that separate from the surface and form in the wake. This distinctive pattern is known as the Von Kármán vortex street, where alternating vortices of a regular pattern are shed downstream of the cylinder. The unsteady Navier-Stokes equations found in equations 14 and 15 are solved using XLB [40], a Lattice Boltzmann method [41] (LBM) library utilizing the JAX framework [42] available for Python, which allows for effective scaling across multiple CPUs, GPUs, and distributed multi-GPU systems. The cited works can be referred to for an overview of the Lattice Boltzmann method and its implementation in XLB.
@@
The diameter $d$ is set to an integer quantity by rounding to the nearest whole number. The wake structure behind the cylinder undergoes instabilities [43, 44] at a critical reynolds number of approximately $Re_{c}=180$, where the vortices transition to becoming turbulent. As a result, the prescribed range given for $Re$ in the parameter space includes a variety of physical regimes. Additionally, both $Re$ and $d$ control the size and periodicity of the shed vortices, making this a difficult prediction problem for ROMs. The freestream velocity in the $x$-direction is set to $u_{\infty}=0.001$ in non-dimensional units. The simulation is run for $T=750000$ timesteps with data being saved every 1000 steps, leading to 750 snapshots for a single simulation. In LBM simulations, time and space are often non-dimensionalized by $\Delta t$ and $\Delta x$ to arrive at dimensionless values of $\Delta x^{*}=\Delta t^{*}=1$. However, one needs to ensure that $\Delta t/\Delta x\ll 1$ to avoid compressibility errors. In these runs we picked $\Delta t/\Delta x=0.001$. The initial condition for a simulation of a given diameter $d$ is the solution to steady flow at $Re=20$.
50 sets of design parameters are generated and split into 45 training samples and 5 test points, shown in Figure 9. Table V also lists the test point design parameters. The ROM uses $m=96$ bagged LSTMs and a window size of $w$ = 30, which were again chosen through a trial-and-error process to maximize accuracy while keeping the number of weak learners to a minimum. The CAE latent dimension is set to $k=10$. Similar to the lid-driven cavity case, values below $k=10$ resulted in higher reconstruction errors and increasing the latent dimension further offers no improvement. A bidirectional LSTM architecture [45] is used. Bidirectional recurrent neural networks process sequential data in both forward and backward directions, allowing the model to learn both past and future context. For problems that are characterized by regularly repeating patterns such as a vortex street, bidirectional LSTMs can better learn the cyclic nature of the pattern by leveraging context in both directions. The network consists of three hidden layers with 36 neurons each and a dropout rate of 0.15. The output layer again contains a sigmoid activation function, and the Adam optimizer with the same initial learning rate $\eta=5\times 10^{-4}$ and weight decay of $\lambda=1\times 10^{-6}$ is used for both the LSTM and CAE. Again, the CAE is trained for 200 epochs while an individual LSTM is trained for 250 epochs. The CAE architecture is given in Appendix B. At the test points, the FOM is run for $T_{i}=300000$ seconds (300 snapshots), or 40% of the total simulation time. This value is required to be high as the flow exhibits highly oscillatory behavior initially, leading to very noisy latent variables that cannot be used for model training. As a result, latent variable sequences are generated starting at the 200th snapshot, and $\bm{A}_{\text{train}}$ does not contain time-series data of latent variables before this point.
@@
Table V: Test case design parameters for the 2D cylinder case.